The brand manager for a brand of toothpaste must plan a campaign designed to increase brand recognition. He wants to first determine the percentage of adults who have heard of the brand. How many adults must he survey in order to be 90​% confident that his estimate is within eight percentage points of the true population​ percentage? Complete parts​ (a) through​ (c) below. b) Assume that a recent survey suggests that about 84​% of adults have heard of the brand.
Answer #1

Solution :

Given that,

р = 0.84

1 - р = 1 - 0.84 = 0.16

margin of error = E = 0.08

At 90% confidence level

\alpha = 1 - 90%

\alpha =1 - 0.90 =0.10

\alpha/2 = 0.05

Z\alpha/2 = 1.645

sample size = n = (Z\alpha / 2 / E )2 * р * (1 - р )

= (1.645 / 0.08)2 * 0.84 * 0.16

= 56.82

sample size = n = 57